OpenStudy (anonymous):

Passes through (3, 2) and is parallel to 8x – 2y = 4. I need the point-slope form and slope intercept form

OpenStudy (anonymous):

@texaschic101

OpenStudy (ivettef365):

First find the slope on the equation 8x - 2y = 4 ??

OpenStudy (anonymous):

so y2-y1/x2-x1?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

no, you use the equation y = mx + b, which m=slope so you have 8x - 2y = 4, isolate the y -2y = -8x + 4 y = 4x + 4 , now the 4 is the m which is the slope

OpenStudy (ivettef365):

now that you know the slope, to be parallel needs to have the same slope so you use the formula again to find b, substitute m = 4, (x,y) for the points (3,2)

OpenStudy (anonymous):

wait so the slope is y=4x+4?

OpenStudy (ivettef365):

to find b, y=mx+b 2 = 4(3) + b, so b = 2-12 = -10 so the equation would be y = 4x -12

OpenStudy (ivettef365):

sorry, I meant y = 4x -10

OpenStudy (anonymous):

going fast, slow down, y=4x-10?

OpenStudy (ivettef365):

just let me know what part you got lost

OpenStudy (anonymous):

the slope =4 or 4x+4

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

the slope = 4

OpenStudy (ivettef365):

y = a point, x = a point an m = slope, b = y-int y = mx + b

OpenStudy (anonymous):

slope intercept form is y=mx+b?

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

that is the formula, where (x,y), could be any points,

OpenStudy (anonymous):

so my slope intercept form equation would be y=4x-10

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

and point slope form is y-y1=m(x-x1) correct?

OpenStudy (ivettef365):

yes, that is the slope intercept form

OpenStudy (ivettef365):

that is correct

OpenStudy (anonymous):

so 4-2=4(x-3)?

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

2=4x-12 y=2x-6 is that my point slope equation? @ivettef365

OpenStudy (ivettef365):

that would be y -2 = 4(x-3)

OpenStudy (anonymous):

got it thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (ivettef365):

yw :)